the area of a triangle is 140 inches squared. If the base and height are both doubled, what is the new area?
Answer by roseo(33)

Remember that the area of a triangle is A = 1/2 bh and your area is 140. Therefore 140 = 1/2 bh. If we double the base and double the height then we have 2b for the base and 2h for the height. Then the area of the new triangle would be 1/2 * 2b*2h or 2bh. If we look at the original triangle with area 140 = 1/2bh then bh = 280 (by multiplying both sides by 2). Now for the new triangle
A = 1/2 bh
= 1/2 (2b)(2h)
=2bh
we know that bh = 280 so 2(280) = 560square inches